OpenStudy (anonymous):

What is the slope of the line for the equation x = 3?

OpenStudy (bahrom7893):

slope is infinity..

OpenStudy (bahrom7893):

or it doesnt exist..

OpenStudy (bahrom7893):

its just a vertical line passing thru 3 on the x axis

OpenStudy (anonymous):

ijust told you in the chat that it was 3 . there is no y intercept so if your graphing the line will be vertical

OpenStudy (sriram):

look slope is tan of the angle btwn the line and x axis x-3 makes 90 degrees wid x axis hence slope= tan 90 = 1/0 hence not defined

OpenStudy (bahrom7893):

It is not 3...it is infinity..equation with slope 3 would be: y=3x+b

OpenStudy (anonymous):

umm excuse me slop is x whats the equation you just gave y=mx+b correct , ok then x=3 which is the slope

OpenStudy (anonymous):

x=slope b=y intercept

OpenStudy (bahrom7893):

okay made for this you are wrong.. sorry but i dont have time to prove it right now..

OpenStudy (anonymous):

ok sure @chiquakia are you doinq a test or somethinq ?

OpenStudy (anonymous):

I know that you guys have been discussing this for a while, but just for the heck of it: Given: x=3 If we transform this into the y=mx+b form, we get: 0y=-x+3 dividing both sides by zero (if we are allowed to) , we get (-1/0) as slope, and (3/0) as the y-intercept. Therefore, the slope is (-1/0) or undefined.

OpenStudy (anonymous):

im help a friend out